How many 6 of diamonds in a deck of 52 cards?
There is only 1 in each deck.
There are 4 '6s' in total:
- 6 of hearts
- 6 of clubs
- 6 of spades
- 6 of diamonds
What is the probability of a number cube?
A number cube has six faces, so the probability of any one of them showing on a fair throw is 1 in 6, or about 0.1667.
68 % is about one standard deviation - so there score would be between 64 and 80 (72 +/- 8)
What is difference between sample and representative sample?
A sample is any subset of the total population. A representative sample is one that is chosen so that its characteristics are similar to that of the population.
Why is heads more likely than tails?
It isnt, There is a 50/50 chance on getting a head or either a tail, as there are 2 sides as it has 2 faces. I hope this helped :D X
How many cards are diamond or a four in a 52 deck of regular cards?
13 cards are diamonds. Four cards are 4. One card is the 4 of diamonds.
What is the probability of drawing a heart and rolling a 4?
The probability of drawing a heart in a single random draw from a normal deck of playing cards and rolling a 4 on a fair number cube is:
1/4*1/6 = 1/24
What is the probability of getting a diamond and a black 7?
The probability of getting a diamond and a black seven is zero. Diamonds are red.
What is the probability of getting the same number on a roll of 3 dice?
(probably but very unlikely)
-----------------------------------
A better answer would be as follows. Throw the first dice. It does not matter which number turns up. Let's suppose it was a 2; Now there is only one 2 out of six different numbers when you throw the second dice. So the probability of scoring another 2 is 1 out of 6 = 1/6 Now for each of the results from the two dice which we have read there is only one 2 from the 6 possible numbers, which again means a probability of 1 out of six = 1/6
So the final probability is found by multiplying 1 (certainty) for the first dice by 1/6 for the second dice by 1/6 for the third dice, which = 1/36
What is the probability of of tossing heads on the first 6 tosses of a fair coin?
The probability of tossing heads on all of the first six tosses of a fair coin is 0.56, or 0.015625.
The probability of tossing heads on at least one of the first six tosses of a fair coin is 1 - 0.56, or 0.984375.
Ap statistics multiple choice preparation book answers?
Reddish peach book (authors: Joyce Heller, Lenore Lerer, and others) answers
Section 1 sample examination 1
1. For the years 1950-1980… A
2. A
3. C
4. D
5. B
6. C
7. B
8. E
9. C
10. E
11. E
12. D
13. B
14. C
15. E
16. D
17. D
18. B
19. B
20. E
21. B
22. C
23. C
24. B
25. D
26. E
27. C
28. B
29. B
30. C
31. D
32. C
33. E
34. D
35. B
36. C
37. D
38. B
39. C
40. A student was interested in the relationship… D
Sample examination 2
1. Creating a sample of students by starting… D
2. D
3. B
4. A
5. B
6. B
7. B
8. B
9. C
10. E
11. E
12. A
13. A
14. D
15. E
16. E
17. D
18. B
19. B
20. D
21. D
22. D
23. A
24. B
25. C
26. B
27. C
28. C
29. E
30. E
31. D
32. B
33. C
34. D
35. D
36. D
37. A
38. C
39. D
40. C
What is the probability of getting a sum greater than 9 when two dice are rolled?
There are 30 possible combinations of die rolls (assuming 2 6-sided die) (excluding "doubles")
The highest total you can get is 12 so the numbers you are after are 10,11 and 12...
The possible combos which add to greater than 9 are:
4,6
6,4
5,5
5,6
6,5
So the probability of rolling a 10, 11 or 12 is 5:30 (5 times out of 30 rolls)
What is one advantage of transforming X values into z-scores?
True or False, One major advantage of transforming X values into z-scores is that the z-scores always form a normal distribution
What are the odds of running out of peanut butter and jelly at the same time?
That's only possible to calculate if more detail is given. For example, how likely is the person to buy jelly normally? How much of it do they consume on average? how likely are they to buy peanut butter? what is their average consumption? Do they actually enjoy eating peanut butter and jelly at the same time? etc...
What are the odds in favor of drawing a card higher than a 9 from an ordinary deck of cards?
If you assume that the Ace is high, then the odds of drawing a card higher than a nine is a standard deck of 52 cards is 20 in 52, or 5 in 13, or about 0.3846.
If you assume that the Ace is low, then the odds of drawing a card higher than a nine is a standard deck of 52 cards is 16 in 52, or 4 in 13, or about 0.3077.
What is the probability that a male will inherit an X-linkedrecessive allele from his father?
There is no possibility that a male will inherit and X-linked recessive allele from his father because for a male child the father only contributes the Y chromosome to his son (of the XY pair he has). If the fater's X chromosome has a recessive allele then it is 100% certain that he will pas this on to all his daughters.
How many standard deviation tend to cover the entire range of scores in a distribution?
Type your answer here...
It depends what percentage of the total data you want to embrace. 99.73% of the total distribution lies between minus to plus 3 standard deviations. That's usually the benchmark range.